OpenStudy (anonymous):

Dorothy is drawing a block on a poster. She knows that the area of the block is x^2 - 7x + 6 square cm. Determine which of the following are possible dimensions of the block, in cm. A. (x – 6)(x – 1) B. (x – 3)(x – 2) C. (x + 6)(x – 1) D. (x + 3)(x – 2)

OpenStudy (anonymous):

@severus4ever @Kkutie7

OpenStudy (anonymous):

A

OpenStudy (anonymous):

-6*(-1)=6 -6+(-1)=-7